An investment was worth $2500 in 1995. By 2000 it had grown to $3500. Let y be the value of the investment in year x, where x = 0 represents 1995. Assuming that the value of the investment increases in a linear fashion, write an equation that models the value of the investment in year x.
Answer by stanbon(75887)

An investment was worth $2500 in 1995. By 2000 it had grown to $3500. Let y be the value of the investment in year x, where x = 0 represents 1995. Assuming that the value of the investment increases in a linear fashion, write an equation that models the value of the investment in year x.
----
You have 2 points: (0,2500) and (5,3500)
------------------
slope = (3500-2500)/(5-0) = 1000/5 = 200
---
intercept = f(0) = 2500
-------
Equation:
f(x) = 200x + 2500
===================
Cheers,
Stan H.
===================